For each of the following tobs values, determine the highest level of significance associated with the decision taken. tobs df Decision

(i) t=4.000 df=17 Reject Ha for a two-tailed test

(ii) t=1.200 df=120 Reject Ha for a one-tailed test

(iii) t=-2.660 df=16 Reject Ha for both one tailed and two-tailed test

(iv) t=-1.586 df=60 Reject Ha for a one-tailed test


1
Expert's answer
2021-06-03T14:10:01-0400

(i) "\alpha=0.0009."

(ii) "\alpha=0.1163."

(iii) one-taled "\alpha=0.0086," two-tailed "\alpha=0.0171."

(iv) "\alpha=0.0590."